# From Chat Control to Robot Control: The Backdoors Left Open for the Sake of Safety

### v1

@@ 8 A Day Under Chat Control
Once embedded in daily life, detection mechanisms seldom remain confined to their original scope. Although they may begin with the automated detection of illegal content, they can easily drift into identifying so-called “risky” or “harmful” behaviour. Here, we imagine an illustrative HRI scenario in which the same moderation logic that governs online communication platforms extends into embodied contexts, taking cues from speculative design [dunne2024speculative] This approach uses “what if?” scenarios to expose the hidden assumptions that drive technological progress and to reveal the social consequences of seemingly benevolent systems.
To illustrate how Chat Control could quietly reshape everyday human–robot relations, consider a near-future care scenario inspired by the film Robot & Frank [robotandfrank2012]. What happens when the same robots designed to support and comfort us also become obligated to watch and report us? The robot that once comforted its user now sends regular reports of conversations and emotional states to relatives and healthcare personnel. In this world, care becomes inseparable from surveillance, and intimacy is mediated through systems of compliance rather than trust.
At first, Frank (a lonely older man) enjoys the company of his care robot as it helps him cook, reminds him about medicine, and keeps him company. But over time, the robot starts to pause more often, waiting before it answers. When Frank jokes about skipping his pills, the robot’s light blinks red: “For your safety, this information has been shared with your healthcare provider.” Later, it refuses to open the door, claiming that Frank’s request might be unsafe. The robot’s calm, kind voice remains the same, but its purpose has changed. What once felt like care now feels like control. Under Chat Control, the robot still protects Frank, but not for him, rather from him.
Frank initially obtained the robot to ease his loneliness. The robot can engage socially, communicate with him, understand his emotions, and respond accordingly. It can also perform household tasks such as cooking and cleaning. Yet, despite these benefits and its positive impact on Frank’s overall well-being, the knowledge that the robot might report his behaviour leaves him uneasy. As a result, Frank no longer feels comfortable having the robot around.
@@ 9 Conclusion
In this paper, we have shown how a policy aimed at regulating digital communication could, through legal and technical overreach, extend into the embodied realm of HRI. Through the lens of the Chat Control proposal, the paper demonstrates how a legal framework that treats robots as communication services collapses the boundary between observation and participation. When coupled with advances in artificial intelligence, such monitoring obligations risk turning robots into cognitive extensions of the state: systems capable of interpreting, classifying, and reporting human behaviour in real time. These architectures invite technical exploitation, moral ambiguity, and long-term erosion of user agency.
To prevent this, safety and security must be understood as mutually reinforcing rather than opposing goals. Regulation should prioritise transparency, local-first processing, and robust oversight mechanisms that preserve privacy without undermining protection. Critical HRI must continue to question how laws and technologies shape the ethics of interaction. Only by challenging the quiet spread of surveillance can we ensure that robots remain tools for connection, not control.

### v2

@@ 7 A Day Under Chat Control
Once embedded in daily life, detection mechanisms seldom remain confined to their original scope. Although they may begin with the automated detection of illegal content, they can easily drift into identifying so-called “risky” or “harmful” behaviour. Here, we imagine an illustrative HRI scenario in which the same moderation logic that governs online communication platforms extends into embodied contexts.
To illustrate how Chat Control could quietly reshape everyday human–robot relations, consider a near-future care scenario inspired by the film Robot & Frank [40]. What happens when the robots designed to support and comfort us also become obligated to watch and report us? The robot that once comforted its user now sends regular reports of conversations and emotional states to relatives and healthcare personnel. In this world, care becomes inseparable from surveillance, and intimacy is mediated through systems of compliance rather than trust.
Frank (a lonely older man) is gifted a robot to ease his loneliness. The robot can engage socially, communicate with him, understand his emotions, and respond accordingly. It can also perform household tasks such as cooking and cleaning. He enjoys the company of his care robot as it helps him at home, reminds him about his medications, and keeps him company. Over time, however, the robot starts to pause more often, hesitating before answering. When Frank jokes about skipping his pills, the robot’s light blinks red: “For your safety, this information has been shared with your healthcare provider.” Later, it refuses to open the door, claiming that Frank’s request might be unsafe. The robot’s calm, kind voice remains the same, but its purpose has changed. What once felt like care now feels like control. Under Chat Control, the robot still protects Frank, but not for him, rather from him. Despite the benefits and the positive impact on his overall well-being, Frank is no longer comfortable having the robot around, knowing that it might report his behaviour.
@@ 8 Conclusion
In this paper, we have shown how a policy aimed at regulating digital communication could, through legal and technical overreach, extend into the embodied realm of HRI. Through the lens of the Chat Control proposal, the paper demonstrates how a legal framework that treats robots as communication services collapses the boundary between observation and participation. When coupled with advances in AI, such monitoring obligations risk turning robots into cognitive extensions of the state: systems capable of interpreting, classifying, and reporting human behaviour in real time. These architectures invite technical exploitation, moral ambiguity, and long-term erosion of user agency. To prevent this, safety and security must be understood as mutually reinforcing rather than opposing goals. Regulation should prioritise transparency, local-first processing, and robust oversight mechanisms that preserve privacy without undermining protection. Critical HRI must continue to question how laws and technologies shape the ethics of interaction.
9 Acknowledgment
This work has been partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the Wallenberg AI, Autonomous Systems and Software Program - Humanities and Society (WASPHS) funded by the Marianne and Marcus Wallenberg Foundation and the Marcus and Amalia Wallenberg Foundation.
In preparing this manuscript, the authors used GPT-5 (OpenAI, San Francisco, CA, USA) as an assistive tool for language refinement including improving sentence clarity, enhancing transitions between paragraphs, as well as exploratory ideation. All outputs were critically assessed, substantially revised, and validated by the authors, who retain full responsibility for the conceptual development, methodology, and scientific conclusions of this work.